The following information is available regarding the total manufacturing overhead of Bursa Mfg. Co. for a recent four-month period.

Machine-
Hours
Manufacturing Overhead
January 5,400 $ 310,000
February 3,200 224,000
March 4,900 263,800
April 2,800 190,000


a-1. Use the high-low method to determine the variable element of manufacturing overhead costs per machine-hour. (Round your answer to 2 decimal places.)

a-2. Use the high-low method to determine the fixed element of monthly overhead cost.

b. Bursa expects machine-hours in May to equal 5,300. Use the cost relationships determined in part a to forecast May's manufacturing overhead costs.

c. Suppose Bursa had used the cost relationships determined in part a to estimate the total manufacturing overhead expected for the months of February and March. By what amounts would Bursa have over- or underestimated these costs?
Answer #1

SOLUTION

A1. Variable cost per machine hour = Change in manufacturing overhead / Change in the level of activity

= (310,000-190,000) / (5,400-2,800)

= 120,000 / 2,600

= $46.15

A2. Fixed costs = Total manufacturing cost - Variable costs

= $310,000 - (5,400*46.15)

= $310,000 - 249,210

= 60,790

B. Total manufacturing costs = Variable cost + Fixed cost

= (5,300 * $46.15) + $60,790

= 244,595 + 60,790

= $305,385

C. February

Variable costs (3,200*46.15) 147,680
Fixed costs 60,790
Estimated manufacturing cost (a) 208,470
Actual manufacturing cost (b) 224,000
Over estimated costs (a-b) 15,530

March

Variable costs (4,900*46.15) 226,135
Fixed costs 60,790
Estimated manufacturing cost (a) 286,925
Actual manufacturing cost (b) 263,800
Under estimated costs (a-b) 23,125